OpenStudy (anonymous):

Answers: 0.0238 4,368 2,380 Some number less than 1

OpenStudy (anonymous):

http://prntscr.com/5phqku

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

why don't u find it by writing what do they equal to? \[\frac{ 16! }{ 5!.11! } = \frac{ 12.13.14.15.16 }{ 5! } = \frac{ 12.13.14.15.16 }{ 120 }\]

OpenStudy (anonymous):

my answer was 4368